OpenStudy (anonymous):

(K^1/3)^-5 OVER (k^8)^1/6= need help with this one.

jimthompson5910 (jim_thompson5910):

1/3 times -5 = ???

OpenStudy (anonymous):

-5/3

jimthompson5910 (jim_thompson5910):

8 times 1/6 = ??

OpenStudy (anonymous):

4/3

jimthompson5910 (jim_thompson5910):

now you subtract the expoennts -5/3 - 4/3 = ???

OpenStudy (anonymous):

gives me -9/3

OpenStudy (anonymous):

well I get a final of 1/k^3

jimthompson5910 (jim_thompson5910):

good, -9/3 = -3

jimthompson5910 (jim_thompson5910):

k^(-3) = 1/( k^3 )

OpenStudy (anonymous):

Thanks Jim. I really appreciate the help :)

jimthompson5910 (jim_thompson5910):

np